OpenStudy (anonymous):

Factor completely x^2 – 12x + 35. (x - 5)(x - 7)?

hero (hero):

yup

hero (hero):

Here are the complete steps: x^2 - 12x + 35 x^2 - 5x - 7x + 35 x(x - 5) -7(x - 5) (x - 5)(x - 7)

OpenStudy (anonymous):

you can use the FOIL method to check this... x-5 x-7 _____ x^2-5x -7x +35 ____________ x^2-12x+35 checks out!!!

hero (hero):

You can also use distributive property to do FOIL as well: (x - 5)(x - 7) = x(x - 5) -7(x - 5) = x^2 - 5x - 7x + 35 = x^2 -12x + 35